# HI-SLAM2: Geometry-Aware Gaussian SLAM for Fast Monocular Scene Reconstruction

### v1

@@ Abstract
We present HI-SLAM2, a geometry-aware Gaussian SLAM system that achieves fast and accurate monocular scene reconstruction using only RGB input. Existing Neural SLAM or 3DGS-based SLAM methods often trade off between rendering quality and geometry accuracy, our research demonstrates that both can be achieved simultaneously with RGB input alone. The key idea of our approach is to enhance the ability for geometry estimation by combining easy-to-obtain monocular priors with learning-based dense SLAM, and then using 3D Gaussian splatting as our core map representation to efficiently model the scene. Upon loop closure, our method ensures on-the-fly global consistency through efficient pose graph bundle adjustment and instant map updates by explicitly deforming the 3D Gaussian units based on anchored keyframe updates. Furthermore, we introduce a grid-based scale alignment strategy to maintain improved scale consistency in prior depths for finer depth details. Through extensive experiments on Replica, ScanNet, and ScanNet++, we demonstrate significant improvements over existing Neural SLAM methods and even surpass RGB-D-based methods in both reconstruction and rendering quality. The project page and source code will be made available at https://hi-slam2.github.io/.
@@ I Introduction
The fundamental challenge in monocular 3D reconstruction stems from the lack of explicit scene geometry measurements [11]. Traditional visual SLAM methods [12, 13, 14, 15, 16, 17] developed over decades but typically provide only sparse or semi-dense map representations, proving insufficient for detailed scene understanding and complete reconstruction. While dense SLAM approaches [18, 19, 20] attempt to address this limitation through per-pixel depth estimation, they remain susceptible to significant depth noise and struggle to achieve complete, accurate reconstructions.
@@
We validate our approach through extensive experiments on both synthetic and real-world datasets, including Replica [37], ScanNet [38], and ScanNet++ [39]. Our method achieves substantial improvements in both reconstruction and rendering quality compared to existing Neural SLAM methods, surpassing even RGB-D-based methods in accuracy. Our method is particularly well-suited for real-time applications that demand rapid and reliable scene reconstruction in scenarios where depth sensors are impractical.
@@ II-B Surface Reconstruction
Surface reconstruction typically follows a two-stage pipeline: camera pose estimation through Structure-from-Motion (SfM) [44, 40], followed by multi-view stereo [45, 46] for dense reconstruction. While widely adopted, these methods are computationally intensive and often produce incomplete reconstructions due to depth estimation uncertainties. Neural implicit representations [47] and their variants [48, 49, 50] have demonstrated high-quality reconstruction capabilities but remain computationally demanding. Recent advances in 3D Gaussian Splatting (3DGS) [51] offer more efficient rendering compared to NeRF-based approaches, and its variants [52, 34] show promising geometry reconstruction capabilities. Our approach leverages 3DGS for efficient scene representation while maintaining high-quality reconstruction, effectively addressing the speed-quality trade-off inherent in previous methods.
@@ II-C Dense Visual SLAM
Figure 2: System Overview: Our framework consists of four key stages: online camera tracking, online loop closing, online mapping, continuous mapping, and offline refinement. The camera tracking is performed using a recurrent-network-based approach to estimate camera poses and generate depth maps from RGB input. For 3D scene representation, we use 3DGS to model scene geometry, enabling efficient online map updates. These updates are integrated with pose graph BA for online loop closing, achieving both fast updates and high-quality rendering. In the offline refinement stage, camera poses and scene geometry undergo full BA, followed by joint optimization of Gaussian primitives and camera poses to further enhance global consistency.
@@ III Methods
Our system is designed to enable fast and accurate camera tracking and scene reconstruction from monocular RGB input. As illustrated in Fig. 2, the system comprises four key components: an online tracker, an online loop closing module, a continuous mapper, and an offline refinement stage. The online camera tracker (Sec. III-A) leverages a learning-based dense SLAM frontend to estimate camera poses and depth maps. Global consistency and real-time performance are achieved through the online loop closure module (Sec. III-B), which combines loop closure detection with efficient Pose Graph Bundle Adjustment (PGBA). For scene representation, we employ 3D Gaussian Splatting (3DGS) (Sec. III-C), enabling efficient online map construction, updates, and high-quality rendering. The offline refinement stage (Sec. III-D) enhances reconstruction quality through full BA and joint optimization of Gaussian map and camera poses ensures optimal global consistency. The final mesh is generated by fusing rendered depth maps through TSDF fusion.
III-A Online Tracking
@@
Our online tracking module builds upon a learning-based dense visual SLAM method [20] to estimate camera pose and depth map of keyframes. By leveraging dense per-pixel information through a recurrent optical flow network, our system can robustly track the camera in challenging scenarios, such as low-textured environments and fast movements. To match per-pixel correspondences among all overlapping frames, we construct a keyframe graph $(\mathcal{V},\mathcal{E})$ which represents the co-visibility relationships between every pair of keyframes. The graph nodes $\mathcal{V}$ correspond to keyframes, each containing a pose $\mathbf{T}\in SE(3)$ and an estimated depth map $\mathbf{d}$. Graph edges $\mathcal{E}$ connect keyframes with sufficient overlap, determined by their optical flow correspondences. To synchronize the estimated states with other modules aiding continuous mapping and online loop closing, a keyframe buffer is maintained to store the information of all keyframes and their respective states.
@@
Incorporate monocular depth prior: To overcome the challenge of depth estimation in difficult areas such as low-textured or occluded regions, we incorporate the easy-to-obtain monocular depth priors [24] into the online tracking process. In the RGB-D mode of [20], depth observations are directly used to compute the mean squared error during BA optimization. However, we can not directly follow this same manner because predicted monocular depth priors exhibit inconsistent scales. To address this, [36] proposes estimating a depth scale and an offset for each depth prior as optimization parameters. Although this approach helps align an overall prior scale, we found that it is not sufficient to fully correct the scale distortions inherent in monocular depth priors.
@@
Matrix $\mathbf{C}$ is diagonal since the depth variables are independent of each other and it can be inverted efficiently as $\mathbf{C}^{-1}=diag(\frac{1}{c_{1}},..,\frac{1}{c_{n}})$ Fig. 3 shows an example of the scale alignment for monocular depth priors. Note that the estimated spatially-varying scales result in well-aligned depth prior with respect to the ground truth depth.
III-B Online Loop Closing
@@
where $\mathbf{H}$ denotes the Hessian matrix. The damping factor $\epsilon=10^{-4}$ and regularization factor $\lambda=10^{-1}$ serve two critical functions: preventing convergence to local minima and improving numerical conditioning, while maintaining rapid convergence. Following optimization, we convert the optimized poses bacl the $SE(3)$ for subsequent tracking. The depth maps are scaled according to the corresponding $Sim(3)$ pose transformations. Additionally, as detailed in Sec. III-C, we update the 3D Gaussian primitives based on the pose updates of their anchor keyframes.
III-C 3D Scene Representation
@@
where $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$ denotes the Gaussian mean and $\mathbf{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ represents the covariance matrix in world coordinates. The covariance matrix $\mathbf{\Sigma}_{i}$ is decomposed into orientation $\mathbf{R}_{i}$ and scale $\mathbf{S}_{i}=diag\{s_{i}\}\in\mathbb{R}^{3\times 3}$, such that $\mathbf{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{T}\mathbf{R}_{i}^{T}$. Each Gaussian also carries attributes for opacity $o_{i}\in[0,1]$ and color $\mathbf{c}_{i}\in\mathbb{R}^{3}$. Unlike the original 3DGS [51], we simplify the color representation by using direct RGB values instead of spherical harmonics, reducing optimization complexity. To handle view-dependent color variations, we employ exposure compensation during the offline refinement stage (Sec. III-D).
@@
where $\mathbf{A}$ denotes the 3×3 color transformation matrix and $\mathbf{b}$ represents the 3×1 bias vector. During the offline refinement stage, these exposure parameters are jointly optimized alongside camera poses and scene geometry, as detailed in Sec. III-D.
Map Management: To ensure that newly observed regions are well represented, we add Gaussians for each new keyframe by unprojecting its estimated depth map into 3D space. To maintain map compactness and prevent redundancy, we apply random downsampling with factor $\psi$ before initializing new Gaussians at the downsampled 3D locations. To control map growth, we implement a pruning strategy that removes Gaussians with low opacity. We reset the opacity values every 500 iterations and perform then interleaved densification and pruning every 150 iterations to balance map size and quality. A detailed analysis of map size evolution is presented in Sec. IV-H.
@@ III-D Offline Refinement
Post-keyframe insertion: The first refinement stage identifies regions with insufficient view coverage, particularly areas near view frustum boundaries. These regions typically arise when forward camera motion is followed by backward rotational movement, as illustrated in Fig. 4. During online processing, keyframe selection relies on average optical flow between neighboring frames, as view coverage cannot be fully evaluated without future trajectory information. To identify under-observed regions in the offline stage, we project each keyframe’s pixels onto its adjacent keyframes and quantify the percentage of pixels that fall outside the fields of view of neighboring keyframes. When this percentage exceeds a predetermined threshold, we flag the region as having insufficient observations. Additional keyframes are then inserted at these locations to achieve more complete scene reconstruction and preserve critical details at scene boundaries.
@@ IV Experiments
To evaluate the performance of the proposed system, we conducted extensive experiments on several challenging datasets, including the synthetic Replica dataset [37] and the real-world datasets ScanNet [38] and ScanNet++ [39]. We begin by providing implementation details and evaluation metrics, followed by quantitative and qualitative comparisons on camera tracking accuracy, geometry, and appearance reconstruction quality against state-of-the-art baselines. Subsequently, we present ablation studies to analyze the impact of different design choices. Finally, we present the runtime performance and map size analysis.
@@ IV-B Datasets
Replica Dataset [37] provides synthetic indoor scenes with high-quality reconstructions, featuring complex geometry and textures. We evaluate using eight RGB-D sequences from [5]. The sequences have perfect camera poses and reconstructions make it ideal for benchmarking dense visual SLAM methods. ScanNet Dataset [38] offers real-world RGB-D captures for 3D scene reconstruction. Following [36], we use eight sequences for tracking evaluation and four additional sequences for geometry reconstruction assessment, using the RGB-D reconstructions as ground truth. ScanNet++ [39] presents a large-scale indoor dataset with laser-scanned ground truth, enabling evaluation of dense SLAM reconstruction quality. While the dataset includes both DSLR and iPhone captures, we specifically evaluate on the iPhone sequences, which present additional challenges due to their lower image quality.
@@ IV-D Camera Tracking Accuracy
We evaluate the camera tracking accuracy of our system against state-of-the-art dense visual SLAM methods on both Replica and ScanNet datasets, including comparisons with RGB-D based approaches. The ATE results in Tables I and II demonstrate the superior tracking accuracy of our system. RGB-D methods SplatTAM [31] and MonoGS [32], despite having access to depth measurements for map-based tracking, achieve lower accuracy than hybrid approaches. Our system, along with Splat-SLAM [33], represents the class of hybrid methods that effectively combine dense SLAM with deep learning foundations. Building upon DROID-SLAM [20], we introduce global pose refinement that greatly enhances tracking accuracy beyond the baseline method.

### v2

@@ Abstract
We present HI-SLAM2, a geometry-aware Gaussian SLAM system that achieves fast and accurate monocular scene reconstruction using only RGB input. Existing Neural SLAM or 3DGS-based SLAM methods often trade off between rendering quality and geometry accuracy, our research demonstrates that both can be achieved simultaneously with RGB input alone. The key idea of our approach is to enhance the ability for geometry estimation by combining easy-to-obtain monocular priors with learning-based dense SLAM, and then using 3D Gaussian splatting as our core map representation to efficiently model the scene. Upon loop closure, our method ensures on-the-fly global consistency through efficient pose graph bundle adjustment and instant map updates by explicitly deforming the 3D Gaussian units based on anchored keyframe updates. Furthermore, we introduce a grid-based scale alignment strategy to maintain improved scale consistency in prior depths for finer depth details. Through extensive experiments on Replica, ScanNet, Waymo Open, ETH3D SLAM and ScanNet++ datasets, we demonstrate significant improvements over existing Neural SLAM methods and even surpass RGB-D-based methods in both reconstruction and rendering quality. The project page and source code will be made available at https://hi-slam2.github.io/.
@@ I Introduction
The fundamental challenge in monocular 3D reconstruction stems from the lack of explicit scene geometry measurements [11]. Traditional visual SLAM methods [12, 13, 14, 15, 16, 17] developed over decades and typically provide only sparse or semi-dense map representations, proving insufficient for detailed scene understanding and complete reconstruction. While dense SLAM approaches [18, 19, 20] attempt to address this limitation through per-pixel depth estimation, they remain susceptible to significant depth noise and struggle to achieve complete, accurate reconstructions.
@@
We validate our approach through extensive experiments on both synthetic and real-world datasets, including Replica [37], ScanNet [38], Waymo Open [39], ETH3D SLAM [40], and ScanNet++ [41]. Our method achieves substantial improvements in both reconstruction and rendering quality compared to existing Neural SLAM methods, surpassing even RGB-D-based methods in accuracy. Our method is particularly well-suited for real-time applications that demand rapid and reliable scene reconstruction in scenarios where depth sensors are impractical.
@@ II-B Surface Reconstruction
Surface reconstruction typically follows a two-stage pipeline: camera pose estimation through Structure-from-Motion (SfM) [46, 42], followed by multi-view stereo [47, 48] for dense reconstruction. While widely adopted, these methods are computationally intensive and often produce incomplete reconstructions due to depth estimation uncertainties. Neural implicit representations [49] and their variants [50, 51, 52] have demonstrated high-quality reconstruction capabilities but remain computationally demanding. Recent advances in 3DGS [53] offer more efficient rendering compared to NeRF-based approaches, and its variants [54, 34] show promising geometry reconstruction capabilities. Our approach leverages 3DGS for efficient scene representation while maintaining high-quality reconstruction, effectively addressing the speed-quality trade-off inherent in previous methods.
@@ II-C Dense Visual SLAM
Figure 2: System Overview: Our framework consists of four key stages: online camera tracking, online loop closing, online mapping, continuous mapping, and offline refinement. The camera tracking is performed using a recurrent-network-based approach to estimate camera poses $\mathbf{T}$ and generate depth maps $\mathbf{D}$ from RGB input. Depth priors are incorporated into the tracking process through our proposed Joint Depth and Scale Alignment (JDSA) strategy improving depth estimation accuracy. For 3D scene representation, we use 3DGS to model scene geometry, enabling efficient online map updates. These updates are integrated with $Sim(3)$-based pose graph Bundle Adjustment (BA) for online loop closing, allowing for scale drift correction via scale updates $\Delta\mathbf{s}$, and achieving both fast updates and high-quality rendering. In the offline refinement stage, camera poses and scene geometry undergo full BA, followed by joint optimization of Gaussian primitives and camera poses to further enhance global consistency.
@@ III Methods
Our system is designed to enable fast and accurate camera tracking and scene reconstruction from monocular RGB input. As illustrated in Fig. 2, the system comprises four key components: an online tracker, an online loop closing module, a continuous mapper, and an offline refinement stage. The online camera tracker (Sec. III-B) leverages a learning-based dense SLAM frontend to estimate camera poses and depth maps. Global consistency and real-time performance are achieved through the online loop closure module (Sec. III-C), which combines loop closure detection with efficient Pose Graph Bundle Adjustment (PGBA). For scene representation, we employ 3DGS (Sec. III-D), enabling efficient online map construction, updates, and high-quality rendering. The offline refinement stage (Sec. III-E) enhances reconstruction quality through full BA and joint optimization of Gaussian map and camera poses ensures optimal global consistency. The final mesh is generated by fusing rendered depth maps through TSDF fusion.
III-A Comparison to HI-SLAM
The current HI-SLAM2 system represents a significant advancement over our previous work HI-SLAM [36], with improvements across multiple aspects that enhance tracking accuracy and reconstruction quality substantially. The key improvements can be summarized as follows:
  • Depth Prior Integration: We propose a novel spatially-adaptive scale-grid alignment strategy that effectively addresses nonlinear scale distortions in monocular depth priors. In contrast to HI-SLAM’s single scale alignment, our 2D grid-based method with bilinear interpolation accommodates spatially varying distortions. This is achieved without introducing dependencies between depth pixels, as each depth value is treated as an independent variable. This design preserves the efficiency of solving the optimization problem using the Schur complement. The improved scale alignment enhances the accuracy of depth estimation, thereby facilitating Gaussian initialization and optimizing the map reconstruction with depth supervision.
  • Map Representation: We replace HI-SLAM’s neural implicit field representation with 3DGS, transitioning from an implicit to explicit representation. This change provides several benefits: (1) significantly faster rendering, (2) efficient online map updates through direct Gaussian primitive updates rather than neural network weight optimization, (3) incremental map growth without predefined scene boundaries, and (4) enhanced geometry preservation as evidenced by the 1.54cm improvement in reconstruction accuracy on the Replica dataset (Table IV).
  • Hierarchical Optimization: Our system employs a multi-stage optimization pipeline that includes online tracking with local BA, online loop closure with $Sim(3)$-based PGBA, and global full BA. Last but not least, the joint refinement of both Gaussian map parameters and camera poses, which couples the tracking frontend with the mapping backend more tightly. In contrast, HI-SLAM only performs pose optimization in the tracking frontend, which can lead to inconsistencies between estimated poses and map geometry. This hierarchical approach reduces the Absolute Trajectory Error (ATE) by 29.3% compared to online tracking alone (Table IX) on the Replica dataset, yielding a globally more consistent reconstruction.
III-B Online Tracking
@@
Our online tracking module builds upon a learning-based dense visual SLAM method [20] to estimate camera poses and depth maps of keyframes. By leveraging dense per-pixel information through a recurrent optical flow network, our system can robustly track the camera in challenging scenarios, such as low-textured environments and fast movements. To match per-pixel correspondences among all overlapping frames, we construct a keyframe graph $(\mathcal{V},\mathcal{E})$ which represents the co-visibility relationships between every pair of keyframes. The graph nodes $\mathcal{V}$ correspond to keyframes, each containing a pose $\mathbf{T}\in SE(3)$ and an estimated depth map $\mathbf{d}$. Graph edges $\mathcal{E}$ connect keyframes with sufficient overlap, determined by their optical flow correspondences. To synchronize the estimated states with other modules aiding continuous mapping and online loop closing, a keyframe buffer is maintained to store the information of all keyframes and their respective states.
@@
Incorporate monocular depth prior: To overcome the challenge of depth estimation in difficult areas such as low-textured or occluded regions, we incorporate the easy-to-obtain monocular depth priors [24] into the online tracking process. In the RGB-D mode of [20], depth observations are directly used to compute the mean squared error during BA optimization. However, we can not directly follow the same manner because predicted monocular depth priors exhibit inconsistent scales. To address this, [36] proposes estimating a depth scale and an offset for each depth prior as optimization parameters. Although this approach helps align an overall prior scale, we found that it is not sufficient to fully correct the scale distortions inherent in monocular depth priors.
@@
The grid resolution is set to $2\times 2$. Higher resolutions may introduce instability, particularly in low-texture regions where optical flow predictions have low confidence, resulting in insufficient weighting in the reprojection error term for each grid tile. The scale coefficients are initially set to ones. After the system converges, the scale coefficients for new depth priors are initialized using those from the previous depth priors.
@@
Matrix $\mathbf{C}$ is diagonal since the scale alignment in Eq. 2 is applied to the depth prior rather than the depth variables. This preserves the independence between the depth variables allowing us to invert $\mathbf{C}$ efficiently as $\mathbf{C}^{-1}=diag(\frac{1}{c_{1}},..,\frac{1}{c_{n}})$. Fig. 3 shows an example of the scale alignment for monocular depth priors. Note that the estimated spatially-varying scales result in well-aligned depth prior with respect to the ground truth depth.
III-C Online Loop Closing
@@
where $\mathbf{H}$ denotes the Hessian matrix. The damping factor $\epsilon=10^{-4}$ and regularization factor $\lambda=10^{-1}$ serve two critical functions: preventing convergence to local minima and improving numerical conditioning, while maintaining rapid convergence. Following optimization, we convert the optimized poses back to the $SE(3)$ for subsequent tracking. The depth maps are scaled according to the corresponding $Sim(3)$ pose transformations. Additionally, as detailed in Sec. III-D, we update the 3D Gaussian primitives based on the pose updates of their anchor keyframes.
III-D 3D Scene Representation
@@
where $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$ denotes the Gaussian mean and $\mathbf{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ represents the covariance matrix in world coordinates. The covariance matrix $\mathbf{\Sigma}_{i}$ is decomposed into orientation $\mathbf{R}_{i}$ and scale $\mathbf{S}_{i}=diag\{s_{i}\}\in\mathbb{R}^{3\times 3}$, such that $\mathbf{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{T}\mathbf{R}_{i}^{T}$. Each Gaussian also carries attributes for opacity $o_{i}\in[0,1]$ and color $\mathbf{c}_{i}\in\mathbb{R}^{3}$. Unlike the original 3DGS [53], we simplify the color representation by using direct RGB values instead of spherical harmonics, reducing optimization complexity. To handle view-dependent color variations, we employ exposure compensation during the offline refinement stage (Sec. III-E).
@@
where $\mathbf{A}$ denotes the 3×3 color transformation matrix and $\mathbf{b}$ represents the 3×1 bias vector. During the offline refinement stage, these exposure parameters are jointly optimized alongside camera poses and scene geometry, as detailed in Sec. III-E.
Map Management: To ensure that newly observed regions are well represented, we initialize a set of 3D Gaussian primitives when each new keyframe is created to populate the Gaussian map. The initialization process begins by unprojecting the estimated depth map of the keyframe into 3D space. Specifically, each pixel’s depth value is back-projected to generate a 3D point, which serves as the mean $\mu_{i}$ of a new Gaussian primitive. The orientation is initialized as the unit orientation. The initial scale $\mathbf{S}_{i}$ is set by finding the average distance of nearest 3 neighbors to adapt to the local point density, while opacity $o_{i}$ is initialized to 0.5 to allow the optimization itself to update. Color attributes $\mathbf{c}_{i}$ are assigned based on the corresponding pixel’s RGB value from the keyframe. To maintain map compactness and prevent redundancy, we apply random downsampling with a factor $\psi$. This downsampling ensures computational efficiency while preserving enough spatial coverage. To control map growth, we implement a pruning strategy that removes Gaussians with low opacity to eliminate redundant or insignificant primitives. We reset the opacity values every 500 iterations and perform interleaved densification and pruning every 150 iterations to balance map size and quality. A detailed analysis of map size evolution is presented in Sec. IV-H.
@@ III-E Offline Refinement
Post-keyframe insertion: The first refinement stage identifies regions with insufficient view coverage, particularly areas near view frustum boundaries. These regions typically arise when forward camera motion is followed by backward rotational movement, as illustrated in Fig. 4. During online processing, keyframe selection relies on average optical flow between neighboring frames, as view coverage cannot be fully evaluated without future trajectory information. To identify under-observed regions in the offline stage, we project each keyframe’s pixels onto its adjacent keyframes and quantify the percentage of pixels that fall outside the fields of view of neighboring keyframes. When this percentage exceeds a predetermined threshold, we flag the region as having insufficient observations. Additional keyframes are then inserted at these locations, and new Gaussian primitives are populated in the same manner as the keyframes inserted during the online process. This ensures more complete scene reconstruction and preserves critical details at scene boundaries.
@@ IV Experiments
To evaluate the performance of the proposed system, we conducted extensive experiments on several challenging datasets, including the synthetic Replica dataset [37] and the real-world datasets ScanNet [38], Waymo [39], ETH3D SLAM [40], and ScanNet++ [41]. We begin by providing implementation details and evaluation metrics, followed by quantitative and qualitative comparisons on camera tracking accuracy, geometry, and appearance reconstruction quality against state-of-the-art baselines. Subsequently, we present ablation studies to analyze the impact of different design choices. Finally, we present the runtime performance and map size analysis.
@@ IV-B Datasets
Replica Dataset [37] provides synthetic indoor scenes with high-quality reconstructions, featuring complex geometry and textures. We evaluate using eight RGB-D sequences from [5]. The sequences have perfect camera poses and reconstructions make it ideal for benchmarking dense visual SLAM methods. ScanNet Dataset [38] offers real-world RGB-D captures for 3D scene reconstruction. Following [36], we use eight sequences for tracking evaluation and six additional sequences for geometry reconstruction assessment, using the RGB-D reconstructions as ground truth. ETH3D SLAM [40] provides a diverse set of real-world RGB-D sequences with motion-capture ground truth poses, featuring challenging conditions such as extreme lighting variations and complete darkness. Waymo Open dataset [39] provides real-world outdoor data with ground truth vehicle poses. The front-view camera images are used as input for our system. Following the evaluation protocol of [66], we evaluate the tracking accuracy and rendering quality using 9 sequences. ScanNet++ [41] presents a large-scale indoor dataset with laser-scanned ground truth, enabling evaluation of dense SLAM reconstruction quality. While the dataset includes both DSLR and iPhone captures, we specifically evaluate on the iPhone sequences, which present additional challenges due to their lower image quality.
@@ IV-D Camera Tracking Accuracy
TABLE III: Camera tracking and rendering results on Waymo open dataset averaged over 9 sequences.
Metrics NICER- SLAM[58] GLORIE- SLAM[59] Photo- SLAM[72] MonoGS[32] OpenGS- SLAM[66]     Ours
ATE [m] $\downarrow$ 19.59 0.536 19.95 8.529 0.839 0.457
PSNR $\uparrow$ 12.22 18.83 17.73 21.80 23.99 28.99
SSIM $\uparrow$ 0.622 0.702 0.741 0.780 0.800 0.872
LPIPS $\downarrow$ 0.726 0.572 0.674 0.577 0.434 0.219
Figure 6: Analysis of successful sequences relative to ATE error thresholds on the ETH3D SLAM dataset.
We evaluate the camera tracking accuracy of our system against state-of-the-art dense visual SLAM methods on indoor Replica, ScanNet, ETH3D SLAM datasets and outdoor Waymo datasets, including comparisons with RGB-D based approaches. The ATE results in Tables I and II demonstrate the superior tracking accuracy of our system. RGB-D methods SplatTAM [31] and MonoGS [32], despite having access to depth measurements for map-based tracking, achieve lower accuracy than hybrid approaches. Our system, along with Splat-SLAM [33], represents the class of hybrid methods that effectively combine dense SLAM with deep learning foundations. The global BA of DROID-SLAM [20] was enabled in all experiments. While it employs global BA, our additional global pose and map joint refinement further improves tracking accuracy beyond the baseline method. We further compare our system on the ETH3D SLAM dataset with state-of-the-art sparse and dense methods, including ORB-SLAM3 [17], DPVO [73], DPV-SLAM [74], and MASt3R-SLAM [75]. Figure 6 illustrates the cumulative success curves based on ATE thresholds. As none of the methods can successfully track all sequences, we use the Area Under the Curve (AUC) metric to assess both accuracy and robustness with an upper ATE threshold of 0.5m. Our system achieves the highest AUC among all methods. Out of total 61 sequences, 6 sequences failed due to complete darkness, while 4 sequences encountered tracking failures caused by lighting changes and view occlusions. These limitations could potentially be addressed in the future by integrating place recognition for relocalization. Our evaluation on the Waymo Open dataset (Table III) further validates our approach, where we achieve the lowest ATE among all competing methods. This highlights the ability of our system to generalize effectively to challenging large-scale outdoor environments with complex scene geometries that typically pose substantial difficulties for monocular systems.